A man buys a certain number of bulbs for ` 300.
He could have bought 5 more bulbs for the same
amount and the bulbs been ` 2 cheaper. Find the
no. of bulbs bought.

Answer by richwmiller(17219)

xy=300,
(x+5)*(y-2)=300
x = 25, y = 12
check
25*12=300
30*10=300
ok